OpenStudy (anonymous):

HELP PLEASE FOR FAN + MEDAL!!! –2x + 4y > 8 Write the equation in point-slope form for the line that is perpendicular to this line and contains the point (1, -4).

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm -2x + 4y > 8 }\) Say it is: \(\normalsize\color{royalblue}{ \rm -2x + 4y = 8 }\) and then please re-write it in a slope intercept form.

OpenStudy (solomonzelman):

Slope intercept form is: \(\normalsize\color{royalblue}{ \rm y=mx+b }\)

OpenStudy (solomonzelman):

Do this for me now, (but it won't be the final answer... beware of that)

OpenStudy (anonymous):

8= -2x+4y is that it @SolomonZelman?

OpenStudy (solomonzelman):

not really.

OpenStudy (solomonzelman):

lets do it together:

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm -2x + 4y = 8 }\) \(\normalsize\color{red}{ \rm +2x~~~~~~~~~~~+2x }\)

OpenStudy (solomonzelman):

what do you then get?

OpenStudy (anonymous):

0 and...10x i think

OpenStudy (solomonzelman):

no,

OpenStudy (solomonzelman):

8 and 2x are not like terms. And on the left side, you left out 4y

OpenStudy (anonymous):

it's 4?

OpenStudy (solomonzelman):

what is 4?

OpenStudy (anonymous):

the left side of the equation -2x + 2 and 4y

OpenStudy (solomonzelman):

\(\normalsize\color{royalblue}{ \rm -2x+4y=8 }\) \(\normalsize\color{royalblue}{ \rm -2x+4y\color{red}{+2x}=8\color{red}{+2x} }\) \(\normalsize\color{royalblue}{ \rm \cancel{-2x}+4y\color{red}{\cancel{+2x}}=8\color{red}{+2x} }\) \(\normalsize\color{royalblue}{ \rm 4y=8+2x }\)

OpenStudy (solomonzelman):

then divide each term in the equation by 4....

OpenStudy (anonymous):

\[y=\frac{ x }{ 2 }+8\] That's what I got is that the equation in slope intercept form @SolomonZelman?

OpenStudy (solomonzelman):

oh wait.

OpenStudy (solomonzelman):

you forgot to divide the 8

OpenStudy (solomonzelman):

\(\large\color{royalblue}{ \displaystyle y=\frac{x}{2}+2 }\)

OpenStudy (anonymous):

oh right hold on let me do it again.

OpenStudy (solomonzelman):

(I divided the 8 too)

OpenStudy (solomonzelman):

I will re-write this, as: \(\large\color{royalblue}{ \displaystyle y=\frac{1}{2}x+2 }\)

OpenStudy (solomonzelman):

and you need the line that is perpendicular to I will re-write this, as: \(\large\color{royalblue}{ \displaystyle y=\frac{1}{2}x+2 }\)

OpenStudy (solomonzelman):

So, for a line with a slope m, a perpendicular line will have a slope of -1/m.

OpenStudy (solomonzelman):

now, if your m is 1/2, then -1/m will be?

OpenStudy (anonymous):

-0.5?

OpenStudy (solomonzelman):

\(\large\color{royalblue}{ \displaystyle -\frac{1}{\frac{1}{2}} }\)

OpenStudy (solomonzelman):

no, you are incorrect-:(`

OpenStudy (anonymous):

hmmm okay hold on

OpenStudy (anonymous):

how do i figure tht out?

OpenStudy (solomonzelman):

\(\large\color{royalblue}{ \displaystyle -\frac{1}{\frac{1}{2}}=-1\div\frac{1}{2}=-1\times 2=?}\)

OpenStudy (anonymous):

-2

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

now, you know that \(\large\color{slate}{ y=2x+b }\) but you need a y-intercept to make the line go through a point (1-4)

OpenStudy (solomonzelman):

I mean point (1,-4) (I left out a comma)

OpenStudy (solomonzelman):

plug in your point and solve for b

OpenStudy (anonymous):

wait i plug in the points 1and -4 into x?

OpenStudy (solomonzelman):

\(\large\color{slate}{ (-4)=2(1)+b }\) b=?

OpenStudy (anonymous):

b=-6?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

So, you have: \(\large\color{slate}{ y=2x-6 }\)

OpenStudy (solomonzelman):

donne

OpenStudy (anonymous):

Thank you so much @SolomonZelman

OpenStudy (solomonzelman):

yw